OpenStudy (anonymous):

Help me please solve integral cos^2 x * sin^2 x dx

hartnn (hartnn):

you can first use the result, \(\large \sin2x = 2 \sin x\cos x\) Square it, what you get ?

OpenStudy (anonymous):

then it will 1/4 integral sin^2 x dx am i right?

OpenStudy (anonymous):

sorry, sin^2 2x dx

hartnn (hartnn):

yes.

hartnn (hartnn):

can you integrate that ?

OpenStudy (anonymous):

and what then?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

i dont know how

hartnn (hartnn):

use the result \(\sin^2 \theta = (1-\cos 2\theta)/2\)

OpenStudy (anonymous):

thanks, i will try

hartnn (hartnn):

good, ask if you don't get it or any other doubts on this.

OpenStudy (anonymous):

thank you, i solved it! :)

hartnn (hartnn):

thats great! welcome ^_^